### v1

# OpenDriveVLA: Towards End-to-end Autonomous Driving with Large Vision Language Action Model

Abstract
We present OpenDriveVLA, a Vision-Language Action (VLA) model designed for end-to-end autonomous driving. OpenDriveVLA builds upon open-source pre-trained large Vision-Language Models (VLMs) to generate reliable driving actions, conditioned on 3D environmental perception, ego vehicle states, and driver commands. To bridge the modality gap between driving visual representations and language embeddings, we propose a hierarchical vision-language alignment process, projecting both 2D and 3D structured visual tokens into a unified semantic space. Besides, OpenDriveVLA models the dynamic relationships between the ego vehicle, surrounding agents, and static road elements through an autoregressive agent-env-ego interaction process, ensuring both spatially and behaviorally informed trajectory planning. Extensive experiments on the nuScenes dataset demonstrate that OpenDriveVLA achieves state-of-the-art results across open-loop trajectory planning and driving-related question-answering tasks. Qualitative analyses further illustrate OpenDriveVLA’s superior capability to follow high-level driving commands and robustly generate trajectories under challenging scenarios, highlighting its potential for next-generation end-to-end autonomous driving. We will release our code to facilitate further research in this domain.
1 Introduction
End-to-end learning frameworks have emerged as a promising paradigm in autonomous driving, enabling perception, prediction, and planning to be jointly optimized within a unified neural network [54, 5]. By leveraging large-scale driving data, these models learn driving policies directly from raw sensor inputs, achieving impressive performance across diverse driving scenarios. Despite notable progress, existing end-to-end driving approaches still face critical challenges, particularly in limited generalization to long-tail scenarios, insufficient understanding of high-level semantics within complex driving scenes, and inflexible reasoning capabilities for task-driven planning [4]. In parallel, large language models (LLMs) and vision-language models (VLMs) demonstrate strong in-context learning, common sense understanding, and zero-short generalization abilities. These emergent capabilities highlight their promise for autonomous driving, especially given the critical need for robust scene understanding of diverse real-world driving conditions [28, 53]. However, directly leveraging existing VLMs for autonomous driving poses fundamental challenges. Firstly, current VLMs are predominantly optimized for static, 2D image-language tasks, leading to poor spatial reasoning performance in dynamic 3D driving environments [34, 50]. Besides, these models frequently produce hallucinated outputs—incorrect or overly confident descriptions—that significantly compromise safety in autonomous driving. Motivated by these limitations, our work answers a central question: How can we harness the emergent capabilities of VLMs to produce reliable driving actions, while mitigating hallucination risks and balancing inference speed and planning effectiveness?
To this end, we propose OpenDriveVLA, a novel vision-language action model designed explicitly for end-to-end autonomous driving. OpenDriveVLA leverages open-source pre-trained language foundation models to generate interpretable and reliable driving trajectories conditioned on multimodal inputs, including 3D environmental perception, ego vehicle state, and driver commands. OpenDriveVLA bridges vision-language understanding and trajectory generation within a unified model. It is rigorously evaluated on open-loop planning and driving-related question-answering benchmarks, achieving leading results and demonstrating superior trajectory generation and driving-scene understanding. Specifically, our contributions are summarized as follows:
  • We present OpenDriveVLA, an end-to-end vision-language action model, which generates reliable driving trajectories conditioned on multimodal inputs.
  • We introduce a hierarchical vision-language feature alignment module, projecting structured 2D and 3D visual tokens into a unified semantic embedding space to facilitate language-guided trajectory generation.
  • We design an agent-env-ego interaction process to capture interactions among the ego vehicle, dynamic agents, and static map elements, significantly enhancing motion forecasting accuracy and trajectory reliability in complex traffic scenarios.
  • Extensive experiments on nuScenes dataset show that OpenDriveVLA establishes new state-of-the-art results in both open-loop planning and driving-related question answering, consistently outperforming prior LLM-based and end-to-end autonomous driving approaches.
2 Related Work
2.1 End-to-End Autonomous Driving
Autonomous driving evolves through two distinct stages. Traditional approaches rely on a modular design, decomposing the system into perception [3, 40, 24], prediction [51, 30, 48], and planning [12, 37] components. While this structure ensures interpretability and allows for independent optimization, they suffer from cascading errors between stages and are not globally optimized for the final planning objective. In contrast, end-to-end autonomous driving frameworks [14, 17, 5] address this by jointly optimizing perception, prediction, and planning within a unified neural network. These models learn driving policies directly from raw sensor inputs, which improves the model’s adaptability to diverse driving conditions. However, existing end-to-end methods still face semantic reasoning bottlenecks—they struggle to fully comprehend high-level scene semantics, infer complex agent interactions, and adapt to dynamic task requirements. Moreover, their decision-making processes remain opaque, making it difficult to diagnose failure cases, especially in long-tail or unseen scenarios.
(a) VLM as additional Caption or QA Head.
(b) VLM as high-level decision-maker.
(c) OpenDriveVLA (Ours): Vision-Language-Action model as end-to-end model via integrated visual perception and language understanding.
Figure 1: Three paradigms of leveraging vision language models for end-to-end autonomous driving.
Figure 2: Illustration of main training stages on OpenDriveVLA, including the hierarchical cross-modality alignment, driving instruction tuning, agent-env-ego interaction, and trajectory planning tuning.
2.2 Multi-modal Large Language Models
Large Language Models demonstrated strong emergent capabilities in in-context learning, instruction following, and reasoning [39, 33]. By training on vast amounts of Internet-scale data, these models acquire extensive world knowledge and exhibit strong adaptability across diverse tasks. Their success has also driven the rise of large VLMs, which extend these capabilities into cross-modal reasoning by integrating vision encoders with language models. State-of-the-art VLMs such as GPT-4V [31], LLaVA [26], DeepSeekVL [41], and Qwen-VL [1] demonstrate strong visual understanding and multimodal reasoning in open-domain tasks. However, these models are primarily trained on static 2D images or videos and exhibit limited spatial reasoning in dynamic 3D driving environments. Moreover, VLMs are prone to hallucinations and generally over-confident but incorrect descriptions, which pose serious risks in safety-critical planning scenarios. Recently, Vision-Language Action models have emerged to directly predict actions from visual inputs, demonstrating strong performance in robotic manipulation tasks [52, 22]. However, the application of such language-conditioned action generation in end-to-end autonomous driving remains underexplored.
2.3 Language Models in Autonomous Driving
Both LLMs and VLMs have been applied to a wide range of autonomous driving tasks, including perception, scene description, synthetic data generation, and high-level decision-making [54]. Within the scope of end-to-end autonomous driving, existing works generally follow three types of design strategies. As shown in Figure 1, one line of research integrates language heads, such as captioning or question-answering modules, into driving models to enhance interpretability [8]. The second category employs language models to generate high-level driving instructions, such as directional commands or abstract maneuvers, which are subsequently interpreted by separate planning modules into low-level controls [18, 38]. This approach allows language models to influence decision-making, but retains a modular separation between reasoning and motion planning, making joint optimization challenging. The third category directly applies VLMs to predict driving actions from monocular video input [19, 44]. These methods process 2D images and generate speed and steering commands based on visual observations, without explicit modeling of the 3D spatial layout or multi-agent interactions in the driving scene. This limits their ability to reason about depth, occlusion, and agent dynamics in complex traffic environments. In this work, we aim to investigate how LLMs can unify cross-modality semantic reasoning and 3D instance-aware trajectory planning within an end-to-end autonomous driving framework.
3 OpenDriveVLA
The overall architecture of OpenDriveVLA is shown in Figure LABEL:fig:title_figure, with its multi-stage training process further detailed in Figure 2. OpenDriveVLA begins with a pre-trained vision encoder, which extracts tokenized environmental representations from multi-view images (Sec. 3.1). These visual tokens are then aligned into the textual domain through cross-modal learning (Sec. 3.2). After alignment, OpenDriveVLA undergoes driving instruction tuning (Sec. 3.3), followed by agent-ego-environment interaction modeling (Sec. 3.4). Finally, OpenDriveVLA is trained end-to-end to predict the ego vehicle’s future trajectory, guided by the aligned visual-language tokens and driving instructions (Sec. 3.5).
3.1 3D Visual Environmental Perception
Recent VLM-based autonomous driving methods typically rely on 2D visual encoders [34, 50], where visual token selection and attention are indirectly guided through language supervision, often in the form of question-answering tasks. While effective in open-domain vision-language applications, this design lacks explicit 3D spatial grounding and structured object-level attention, which can lead to severe hallucinations in safety-critical driving scenarios [42]. To mitigate this, OpenDriveVLA adopts a visual-centric query module, where the model first learns to focus on driving-relevant objects and map tokens through 3D vision tasks, ensuring reliable visual token proposal.
Specifically, given a set of multi-view images ${I}=\{I^{i}\}_{i=1}^{N}$, the visual module first extracts multi-scale 2D features from each image using a shared 2D backbone, denoted as $f_{2D}$. These 2D features are then aggregated across views and lifted into BEV space, producing the BEV feature $f_{bev}$. To obtain structured environmental representations, we adopt three visual query modules: Global Scene Sampler $\mathcal{Q}_{\mathrm{scene}}$, Agent QueryTransformer $\mathcal{Q}_{\mathrm{agent}}$, and Map QueryTransformer $\mathcal{Q}_{\mathrm{map}}$. Each module extracts tokens focusing on a specific semantic aspect of the driving environment. Global Scene Sampler encodes the surrounding driving scene context from multi-view 2D features, producing the scene token $v_{scene}=\mathcal{Q}_{\mathrm{scene}}(f_{2D})$. Agent QueryTransformer detects and tracks dynamic agents within the scene, extracting agent-centric tokens $\{v_{agent}^{i}\}_{i=1}^{N_{a}}=\mathcal{Q}_{\mathrm{agent}}(f_{bev})$, where $N_{a}$ denotes the number of detected agents. In parallel, Map QueryTransformer extracts static structural information, such as lane boundaries and drivable areas, forming the map token $v_{map}=\mathcal{Q}_{\mathrm{map}}(f_{bev})$. Through vision-centric perception tasks, including 3D detection, tracking, and segmentation, the visual encoder produces structured environmental tokens that capture both dynamic agent behaviors and static map structures in a spatially grounded manner. The output tokens, denoted as $\textbf{V}_{env}=\{v_{scene},v_{agent},v_{map}\}$, serve as visual environment representation of the subsequent stages.
3.2 Hierarchical Vision-Language Alignment
To bridge the modality gap between the extracted visual tokens and the word embedding space of a pre-trained LLM, we adopt a hierarchical vision-language feature alignment strategy. Given the visual tokens extracted from the 3D visual perception module, we introduce three token-specific projectors $\{\Phi_{\mathrm{scene}},\Phi_{\mathrm{agent}},\Phi_{\mathrm{map}}\}$. During training, each active agent query from the 3D detection and tracking task denoted as $v_{agent}^{i}$, is also matched to its corresponding ground-truth caption $\mathbf{X}_{agent}^{i}$. These captions provide detailed descriptions, including 2D appearance descriptions and 3D spatial positions. For scene and map tokens, which encode holistic spatial context and static structural properties, a sample-wise alignment is applied, where each token is matched to a scene-level caption $\mathbf{X}_{scene}$ or $\mathbf{X}_{map}$. The scene token $v_{scene}$ captures the global 2D environmental context, while the map token $v_{map}$ encodes structural elements such as lane topology, road boundaries, and drivable areas. Each of these tokens is aligned to its corresponding caption, denoted as $\mathbf{X}_{scene}$ and $\mathbf{X}_{map}$. During this stage, both the visual encoder and the pre-trained LLM remain frozen, with only the token-specific projectors being trainable. The forward alignment step is formulated as follows:
$\hat{\mathbf{X}}_{k}=\mathrm{LLM}\left(\Phi_{k}(v_{k})\right),\quad k\in\{\mathrm{scene},\mathrm{map}\}$ (1)
$\hat{\mathbf{X}}_{agent}^{i}=\mathrm{LLM}\left(\Phi_{\mathrm{agent}}(v_{agent}^{i})\right),\quad i=1,\dots,N_{a}$ (2)
3.3 Driving Instruction Tuning
To embed high-level driving knowledge into OpenDriveVLA and enhance its reasoning ability, we introduce a dedicated driving instruction tuning stage during training. Instead of performing explicit chain-of-thought (CoT) reasoning at inference time, which would significantly increase latency, we distill essential driving knowledge into the model through supervised instruction tuning to achieve a trade-off between inference speed and reasoning efficacy.
During the tuning process, driving knowledge from the language domain is injected into the model using a curated driving instruction QA dataset. The dataset covers a wide range of driving-related reasoning, including perception understanding, motion prediction, attention allocation, action reasoning, and high-level decision-making. By training on this diverse set of driving queries, OpenDriveVLA learns to contextualize the driving scene, follow commands, and generate semantically and behaviorally grounded planning decisions. We formulate the tuning data as instruction-response pairs $\{\mathbf{X}_{input},\mathbf{X}_{answer}\}$, where $\mathbf{X}_{input}=(\textbf{V}_{env},\textbf{S}_{ego},\mathbf{X}_{query})$. Here, $\mathbf{X}_{query}$ denotes the driving-related question, and $\textbf{S}_{ego}$ encodes the textual ego vehicle state. Given this multimodal input, the LLM autoregressively learns to generate the target response. During instruction tuning, the visual encoder remains frozen while the token-specific projectors and the LLM are set to be trainable. The instruction prediction process is as:
$\hat{\mathbf{X}}_{answer}=\mathrm{LLM}\left(\textbf{V}_{env},\textbf{S}_{ego},\mathbf{X}_{query}\right)$ (3)
3.4 Agent Environment Ego Interaction
Reliable trajectory planning in autonomous driving necessitates a spatially grounded 3D representation of the environment. Beyond perception, it must also understand dynamic interactions between the ego vehicle and surrounding agents. Effective interaction modeling is essential to ensure that planned trajectories are both feasible and collision-free under real-world driving constraints. However, existing pre-trained LLMs lack an inherent inductive bias for spatial reasoning in 3D driving scenes, as they are predominantly trained on 2D vision-language and text-based datasets. To address the limitation, we introduce a conditional agent motion forecasting task, which serves as a proxy task for 3D agent-environment-ego interaction modeling, allowing the model to learn spatially grounded motion patterns. During this stage, OpenDriveVLA captures the underlying structure of multi-agent dynamics, enhancing its capability for scene-aware trajectory generation and improving decision-making in complex traffic scenarios.
Given scene and map tokens, as well as the ego vehicle state $\textbf{S}_{ego}$, the LLM predicts the future motion of each detected agent based on its projected visual embedding $\Phi_{\mathrm{agent}}(v_{agent}^{i})$. The future motion of agent $a_{i}$ is represented as a sequence of waypoints $\mathcal{W}^{i}_{a}$. The predicted trajectory is conditioned on the scene context, map structure, and ego vehicle state, enabling OpenDriveVLA to infer interaction-aware and spatially grounded motion sequences. The learning objective for the $i$-th agent is formulated as:
$\operatorname*{max}\prod_{t=1}^{T}p\left(w^{i}_{t}\mid w^{i}_{1:t-1},\textbf{V}_{env},\textbf{S}_{ego}\right)$ (4)
This provides OpenDriveVLA with essential spatial priors, enabling it to bridge the gap between high-level semantic reasoning and physically grounded motion planning.
3.5 End-to-end Trajectory Planning
In this stage, OpenDriveVLA plans future driving actions as a sequence of waypoints over the next few seconds, denoted as $\mathcal{W}_{ego}=\{w_{1},w_{2},\dots,w_{T}\}$. Each waypoint $w_{t}$ represents the 2D coordinates $(x_{t},y_{t})$ of the ego vehicle at time step $t$. To enable autoregressive generation using the LLM, the waypoints are first tokenized into a sequence of discrete textual tokens: $\mathcal{T}{traj}=\mathrm{Tokenizer}(\mathcal{W}_{ego})$. The generation process is then cast as a causal sequence prediction task, where each token is predicted conditioned on the visual perception tokens $\textbf{V}_{env}$, the ego state $\textbf{S}_{ego}$, and the driving command $\textbf{X}_{dri}$:
$\hat{\mathcal{T}}_{traj}=\operatorname*{argmax}_{\textbf{T}_{traj}}\prod_{t=1}^{T}p\left(w_{t}\mid w_{1:t-1},\textbf{V}_{env},\textbf{S}_{ego},\textbf{X}_{dri}\right)$ (5)
The whole model, including the visual encoder, cross-modal projectors, and the LLM, is jointly optimized end-to-end during training. At inference, the model autoregressively generates the tokenized trajectory $\hat{\mathcal{T}}_{traj}$, which is then decoded back into numerical waypoints:
$\hat{\mathcal{W}}_{ego}=\mathrm{Decoder}(\hat{\mathcal{T}}_{traj})$ (6)
4 Experiments
4.1 Training Data
We conduct experiments on nuScenes dataset [2], following standard data split into training and validation sets. OpenDriveVLA is trained using the training set paired with corresponding QA captions, while the validation set is exclusively used for performance evaluation to ensure fair comparisons with prior works. We show the training data information for each stage in Table 1.
Vision-Language Alignment. For agent feature alignment, we post-process instance captions from [20], which provides the 2D visual description of individual objects. To further enhance spatial awareness, each object caption is augmented with its corresponding BEV coordinates, enabling the model to associate object attributes with precise spatial locations. For scene tokens, we process multi-view scene descriptions from [46], merging them into unified summaries that describe the driving environment across all camera views. For map tokens, structured language descriptions are derived from ground-truth annotations, translating map elements such as lane dividers, crosswalks, and road boundaries into descriptive text.
Driving Instruction Tuning. We adopt multiple instruction-oriented datasets derived from nuScenes to incorporate driving-specific knowledge into OpenDriveVLA. We unify several datasets into a standardized instruction-based QA format, including driving-related question-answer pairs collected from nuCaption [46], nuScenesQA [32], and nuX [8] dataset. Each QA pair is conditioned on structured environmental visual tokens and the ego vehicle state, ensuring consistency across different data sources. This multimodal instruction tuning process allows OpenDriveVLA to effectively ground language understanding into both environmental perception and scene understanding, bridging perception, reasoning, and action within the language space.
Stage 1 Stage 2 Stage 2.5 Stage 3
Data Annos 2D/3D Cap Instruct QA Trajectory Trajectory
#Samples 536k 566k 459k 28k
Model Trainable Proj Proj, LLM Proj, LLM Full Model
0.5B Size (MB) 3.1 496.9 496.9 552.6
3B Size (MB) 14.2 3099.6 3099.6 3155.3
7B Size (MB) 41.3 7654.2 7654.2 7709.9
Train 0.5B LR ($\psi_{\text{vision}}$) - - - $1\times 10^{-5}$
0.5B LR ($\{\theta_{\text{proj/LLM}}\}$) $1\times 10^{-4}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$
Epoch 1 1 1 1
Table 1: Overview of the multi-stage training pipeline in OpenDriveVLA, including data, model, and training information.
Method ST-P3 metrics UniAD metrics LLM Input
L2 (m) ↓ Collision (%) ↓ L2 (m) ↓ Collision (%) ↓
1s 2s 3s Avg. 1s 2s 3s Avg. 1s 2s 3s Avg. 1s 2s 3s Avg.
None-Autoregressive Methods
ST-P3 [13] 1.33 2.11 2.90 2.11 0.23 0.62 1.27 0.71 - - - - - - - - - Visual
VAD [17] 0.17 0.34 0.60 0.37 0.07 0.10 0.24 0.14 - - - - - - - - - Visual
Ego-MLP [49] 0.46 0.76 1.12 0.78 0.21 0.35 0.58 0.38 - - - - - - - - - Ego
UniAD [14] 0.44 0.67 0.96 0.69 0.04 0.08 0.23 0.12 0.48 0.96 1.65 1.03 0.05 0.17 0.71 0.31 - Visual
InsightDrive [35] 0.23 0.41 0.68 0.44 0.09 0.10 0.27 0.15 0.30 0.72 1.41 0.81 0.08 0.15 0.84 0.36 - Visual
FF [11] - - - - - - - - 0.55 1.20 2.54 1.43 0.06 0.17 1.07 0.43 - LiDAR
EO [21] - - - - - - - - 0.67 1.36 2.78 1.60 0.04 0.09 0.88 0.33 - LiDAR
Autoregressive Methods
DriveVLM [38] 0.18 0.34 0.68 0.40 0.10 0.22 0.45 0.27 - - - - - - - - Qwen-VL-7B Visual
GPT-Driver [29] 0.20 0.40 0.70 0.44 0.04 0.12 0.36 0.17 0.27 0.74 1.52 0.84 0.07 0.15 1.10 0.44 GPT-3.5 Textual
RDA-Driver [15] 0.17 0.37 0.69 0.40 0.01 0.05 0.26 0.10 0.23 0.73 1.54 0.80 0.00 0.13 0.83 0.32 LLaVa-7B Visual
EMMA [16] 0.14 0.29 0.54 0.32 - - - - - - - - - - - - Gemini [36] Visual
OpenEMMA [43] 1.45 3.21 3.76 2.81 - - - - - - - - - - - - Qwen-VL-7B Visual
DME-Driver [10] - - - - - - - - 0.45 0.91 1.58 0.98 0.05 0.28 0.55 0.29 LLaVa-7B Visual
OpenDriveVLA-0.5B (Ours) 0.15 0.32 0.57 0.35 0.01 0.06 0.20 0.09 0.21 0.60 1.22 0.68 0.00 0.15 0.63 0.26 Qwen2.5-0.5B Visual
OpenDriveVLA-3B (Ours) 0.14 0.30 0.55 0.33 0.02 0.07 0.22 0.10 0.19 0.58 1.24 0.67 0.02 0.18 0.70 0.30 Qwen2.5-3B Visual
OpenDriveVLA-7B (Ours) 0.15 0.31 0.55 0.33 0.01 0.08 0.21 0.10 0.20 0.58 1.21 0.66 0.00 0.22 0.55 0.25 Qwen2.5-7B Visual
Table 2: Open-Loop planning performance comparison of different driving models, including both autoregressive methods and non-autoregressive methods. OpenDriveVLA shows powerful planning ability and achieves best-in-class results among open-source models, even with the 0.5B version. We refer to the result summary from [35, 29, 23, 15].
Motion Forecasting and Trajectory Prediction. We formulate both agent motion forecasting and ego trajectory planning in the ego vehicle system, where the model directly predicts future displacements within each entity’s local coordinate frame relative to the ego vehicle for planning and relative to each agent for forecasting. This formulation captures motion dynamics in a spatially consistent manner across all entities. Following [29], the ego vehicle state is encoded as textual input to ensure ego awareness throughout the training process. Both tasks predict 3-second future trajectories, sampled at 0.5-second intervals, resulting in 6 waypoints per trajectory.
4.2 Evaluations
We assess OpenDriveVLA on open-loop planning task on nuScenes benchmark, where the model is evaluated following both ST-P3 [13] and UniAD [14] settings. The evaluation metrics include L2 displacement errors at 1, 2, and 3 seconds, along with the average collision rate over the prediction horizon. To evaluate the scene understanding ability of OpenDriveVLA, we assess its performance on three driving VQA datasets directly after the driving instruction tuning stage, i.e., [46], nuScenesQA [32], and nuX [8]. The VQA evaluation adopts standard MLG metrics, including BLEU, METEOR, CIDEr, BERT-Score, etc.
4.3 Implementation Details
The 3D visual perception module in OpenDriveVLA follows the vision-centric design from [14], using a ResNet-101 backbone for 2D feature extraction. The perception backbone is pre-trained via multi-task learning on 3D object detection, object tracking, and map segmentation. The resulting BEV feature map has a spatial resolution of $200\times 200$. To construct a unified scene representation, the global SceneSampler applies 2D adaptive pooling to each camera view, subsequently concatenating the pooled multi-view features into a global scene token. Agent and map tokens are extracted from the final layer of their respective QueryTransformer modules. Each token type is then mapped into the language space using a separate two-layer MLP with GeLU activation. We adopt Qwen 2.5-Instruct [45] as the pre-trained LLM, which undergoes full parameter tuning during training. The model is trained on 4 NVIDIA H100 GPUs with a batch size of 1, completed in approximately two days. During inference, we set the decoding temperature to 0 to ensure deterministic trajectory generation. We freeze the 2D backbone during stage 3. Detailed training configurations are summarized in Table 1.
Models CIDER BLEU-4 METEOR ROUGE-L
Hint-UniAD [8] 21.7 4.2 12.7 27.0
Hint-VAD [8] 22.4 4.2 13.2 27.6
GPT-4o [44] 19.0 4.0 10.3 24.9
Gemini 1.5 [36] 17.6 3.4 9.3 23.4
Vote2Cap-DETR [6] 15.3 2.6 10.9 24.2
TOD3Cap [20] 14.5 2.5 10.5 23.5
OpenDriveVLA
0.5B (Ours) 32.3 5.4 12.5 27.9
3B (Ours) 25.5 4.3 12.8 27.8
7B (Ours) 26.2 4.5 12.8 27.4
Table 3: Performance comparison on the Nu-X dataset [8]. OpenDriveVLA outperforms prior works in most metrics.
Method nu-Caption nuScenes-QA
BL-1 BL-2 BL-3 BL-4 BERT-S Ext Cnt Obj Sts Cmp H0 H1 Acc
Mini-GPT4 [55] 15.0 6.8 3.7 2.6 84.4 - - - - - - - -
Instruct-BLIP [7] 18.7 13.4 7.4 5.2 85.9 - - - - - - - -
LLaMA-AdapV2 [9] 30.2 17.3 10.4 7.5 86.5 19.3 2.7 7.6 10.8 1.6 15.1 4.8 9.6
LLaVA1.5 [27] 20.0 12.1 8.6 5.4 85.0 45.8 7.7 7.8 9.0 52.1 25.7 41.5 26.2
LiDAR-LLM [47] 41.0 30.0 23.4 19.3 91.3 74.5 15.0 37.8 45.9 57.8 - - 48.6
BEVDet+BUTD [32] - - - - - 83.7 20.9 48.8 52.0 67.7 - - 57.0
OpenDriveVLA-0.5B (Ours) 47.2 35.8 29.4 25.2 91.9 83.9 22.0 50.2 57.0 68.4 62.3 56.5 58.4
OpenDriveVLA-3B (Ours) 48.3 36.9 30.3 26.1 92.0 84.0 22.3 50.3 56.9 68.5 62.6 56.5 58.5
OpenDriveVLA-7B (Ours) 49.6 38.3 31.9 27.6 92.2 84.2 22.7 49.6 54.5 68.8 62.4 56.1 58.2
Table 4: Comprehensive performance comparison across nu-Caption [46] and nuScenes-QA [32]. Here, BL-1/2/3/4 refer to BLEU scores at different n-gram levels, BERT-S refers to BERT-Score, and metrics under nuScenes-QA correspond to different question categories: Ext (Existence), Cnt (Counting), Obj (Object), Sts (Status), and Cmp (Comparison), with Acc representing overall accuracy.
4.4 Main Results
Open Loop Trajectory Planning. We evaluate OpenDriveVLA on the open-loop trajectory planning task using both ST-P3 and UniAD metrics, ensuring comprehensive performance assessment across spatial accuracy and collision avoidance. As shown in Table 2, OpenDriveVLA achieves state-of-the-art performance across both settings. Specifically, both 3B and 7B version OpenDriveVLA achieve an average L2 error of 0.33m under ST-P3 metrics, outperforming prior autoregressive language models [29, 38]. On the UniAD metrics, OpenDriveVLA-7B also achieves great performance with an average L2 error of 0.66m. These results validate the effectiveness of OpenDriveVLA. Notably, OpenDriveVLA-0.5B, despite having significantly fewer parameters, also achieves competitive performance and outperforms prior larger models by a substantial margin. This highlights the efficiency of OpenDriveVLA, enabling strong spatial and semantic reasoning even with reduced model size, making it an effective and scalable solution for language-guided planning.
Driving Question Answering. We evaluate OpenDriveVLA on the driving VQA task across three nuScenes-based datasets, as shown in Table 4 and Table 3. OpenDriveVLA reaches best-in-class performance across all three datasets, consistently outperforming previous language-enhanced driving models and general-purpose multimodal baselines among most metrics. On nuCaption dataset, OpenDriveVLA achieves the best captioning performance among all evaluated models, outperforming both general multimodal LLMs LLaVA1.5 [27] and Mini-GPT4 [55], as well as autonomous driving-specific models such as LiDAR-LLM [47]. For nuScenesQA dataset, OpenDriveVLA also achieves strong performance. Compared to models that directly fuse BEV features with language models such as BEVDet+BUTD [32], OpenDriveVLA demonstrates clear advantages in object and status-related questions, which highlights the benefit of its spatially grounded visual-language alignment. Notably, OpenDriveVLA-0.5B outperforms even the larger 7B on the Nu-X dataset, which shows its powerful scene-understanding ability even with lightweight LLMs.
4.5 Ablation Study
Our ablation study examines the influence of input modalities on OpenDriveVLA’s trajectory planning. The results in Table 5 show that visual inputs enhance the action-making process, while textual high-level commands and historical state information further refine trajectory generation, indicating the contribution of semantic intent and temporal context. Besides, we find that ego-state information plays a significant role in trajectory prediction on nuScenes open-planning benchmark, aligning with prior findings [25]. Due to the imbalanced distribution of the dataset, where many scenarios involve maintaining the current state, the model tends to over-rely on ego-state history, leading to a bias toward conservative predictions. Hence, to further evaluate the generalization and instructed action-following capability of OpenDriveVLA, we conduct qualitative studies to examine its ability to execute diverse actions in response to different human commands.
Visual Ego Hist Cmd Avg. L2 (m) ↓ Avg. Collision (%) ↓
UniAD ST-P3 UniAD ST-P3
1.34 0.75 0.77 0.24
1.30 0.75 1.14 0.49
0.77 0.39 0.29 0.10
0.80 0.40 0.33 0.13
0.68 0.35 0.26 0.09
Table 5: Ablation study on the effect with different input combinations on OpenDriveVLA-0.5B model. The table shows the results under UniAD and ST-P3 metrics.
Figure 3: Visualization of OpenDriveVLA-7B planning actions under different driving instructions and the visual question answering results. The agent motion prediction results are visualized after the agent-env-ego interaction stage. QA showcases are from results in Table 3, 4.
Figure 4: Qualitative results comparison of open-loop planning. The top figure shows the planning results of UniAD [14], while the bottom figure presents those of OpenDriveVLA-7B (Ours). The agent motion prediction results are visualized after the agent-env-ego interaction stage.
4.6 Qualitative Results
Driving Command Following. Figure 3 illustrates trajectory planning at an intersection under two different driver instructions: keep forward and turn right, with the right turn as the ground truth. The results demonstrate OpenDriveVLA’s ability to accurately interpret driver commands and generate corresponding trajectory actions. The predicted trajectories adhere to the given command while remaining environment-aware and contextually reasonable, demonstrating strong command-following ability and generalization capability of our model in complex driving scenes.
Comparison with Prior Methods. Figure 4 compares the open-loop planning results of OpenDriveVLA-7B with UniAD [14]. The qualitative results demonstrate that OpenDriveVLA generates more stable and adaptable trajectories, as UniAD tends to overreact to parked vehicles on the right in narrow-road scenarios. Compared to UniAD, OpenDriveVLA effectively maintains trajectory smoothness and environmental awareness, demonstrating its improved capability for handling complex driving scenarios.
4.7 Discussions and Limitation
While OpenDriveVLA performs well across benchmarks, some limitations remain. It lacks explicit CoT reasoning, relying instead on implicit reasoning from driving instruction tuning, which may weaken inference capabilities in complex scenarios. Besides, despite limited token inputs and outputs, its autoregressive nature hampers real-time deployment in high-speed driving, requiring further optimization. Additionally, its evaluation in an open-loop setting does not account for robustness in interactive traffic environments.
5 Conclusion
In this work, we present OpenDriveVLA, a scalable vision-language action model designed for end-to-end autonomous driving. By leveraging pre-trained large language models, OpenDriveVLA generates reliable driving actions. We propose a hierarchical vision-language feature alignment process and an agent-env-ego interaction mechanism to enhance spatial reasoning and trajectory prediction. Through a multi-stage training paradigm, OpenDriveVLA achieves state-of-the-art performance in open-loop planning and driving-related question answering. Extensive evaluations on nuScenes dataset demonstrate its superior trajectory planning capabilities compared to existing auto-regressive and end-to-end driving models. Our work demonstrates the feasibility of a scalable vision-language-driven approach for autonomous driving and highlights the potential of large language models as a foundation for end-to-end driving action systems.